OpenStudy (anonymous):

for sequence of 3/16,-4/25,5/36,-6/49

Parth (parthkohli):

\[\dfrac{3}{4^2},\dfrac{\color{#C00}-4}{5^2}, \dfrac{5}{6^2},\dfrac{\color{#C00}- 6}{7^2},\underline{}?\]

Parth (parthkohli):

Does that give you a hint?